# A monolithic first–order BSSNOK formulation of the Einstein–Euler equations and its solution with path-conservative finite difference CWENO schemes

### v1

@@ IV.1 Linearized gravitational wave
Figure 1: Linearized gravitational wave test solved with different version of the CWENO scheme. Left panel: $\tilde{A}_{zz}$ component of the extrinsic curvature at the final time, compared to the exact solution. Right panel: time evolution of the Einstein constraints.
@@ IV.2 The robust stability test
Fig. 2 shows the results of our calculations, by reporting the evolution of the Einstein constraints in a battery of tests where the CWENO5 version of the scheme is adopted. These numbers and the corresponding results are certainly satisfactory, even if not of the same quality as those reported in Fig. 4 of [27], where a DG scheme was used in combination with the Z4 formulation of the Einstein equations  [94, 22].
@@ IV.3 The gauge wave
We have first run this test with a small wave amplitude $A=0.01$ over a rectangular domain of size $\Omega=[-0.5,0.5]\times[-0.02,0.02]$ using a CWENO7 scheme. The grid is composed of $256\times 4$ elements, uniformly distributed, and the final time is $t=1000$. As expected, the gauge wave test manifests several pathologies when evolved within the BSSNOK formulation over a long timescale. We have confirmed this behaviour, which is clearly in-printed in the exponential growth of the Einstein constraints, as reported in the left panel of Figure 3. In the right panel we can appreciate the mismatch of the final solution at $t=1000$ with respect to the initial conditions, taking the metric function $\tilde{\gamma}_{xx}$ as a representative quantity.
In spite of these deficiencies, it is still possible to perform a convergence analysis upon this solution, provided the final time is short enough in such a way that the pathologies highlighted above do not have the time to spoil the numerical properties of the scheme. Tab. 2 contains the result of this analysis for a gauge wave with $A=0.1$ at the final time $t=1$ and it essentially confirms the nominal orders of convergence of the scheme.
@@ IV.4 Special relativistic Riemann Problems in the Cowling approximation
All together, these calculation prove the ability of the numerical scheme in treating strong discontinuities in the relativistic regime. At this point we would like to emphasize again that the employed conservative CWENO finite difference scheme is exactly the same as used previously for the evolution of vacuum spacetimes.
@@ IV.5 Michel accretion
We have solved this test over a computational domain $(r,\theta)\in[0.5;10]\times[0+\epsilon;\pi-\epsilon]$, with $\epsilon=0.05$ and covered by a $128\times 16$ uniform grid. The relevant parameters are chosen as $r_{c}=5$ and $\rho_{c}=1.006\times 10^{-7}$, while an ideal gas with adiabatic index $\gamma=5/3$ is assumed. In addition, the initial condition has been perturbed in the rest mass density using a Gaussian profile peaked at the critical radius, with a perturbation amplitude given by $\delta\rho=10^{-1}\rho_{c}$. The well–balancing property described in Sect. III.2 and proposed in [27] has been activated, in combination with the CWENO5 version of the scheme. In spite of the perturbation, and by virtue of the well–balancing, the solution relaxes back to perfect equilibrium. In Fig. 7 we have reported the final profiles of the mass density and of the radial velocity at $t=1000$.
@@
Figure 7: Spherical accretion of matter onto a non-rotating black hole using Kerr-Schild coordinates. The initial condition is compared to the solution at time $t=1000\,M$, both for the rest mass density (left panel) and for the radial velocity (right panel). A CWENO5 scheme with Well Balancing is adopted.
IV.6 The equilibrium TOV star
Another fundamental test in numerical general relativity consists of keeping an equilibrium star model stationary over long timescales, providing the first case in our sample of tests where the full Einstein+Euler equations are addressed. Hence, we have first solved the standard Tolman–Oppenheimer–Volkoff (TOV) system [101, 102, 75, 103] for a simple polytropic gas obeying $p=K\,\rho^{\gamma}$.
After choosing the parameters as in [104], namely a central rest mass density $\rho_{c}=1.28\times 10^{-3}$, $K=100$ and $\gamma=2$, we have integrated the ODEs of the TOV system by a Discontinuous Galerkin solver like in [27], obtaining a total mass $M=1.4\,M_{\odot}$ and a radius $R=14.15\,{\textrm{km}}$. A small perturbation in the pressure is then added to the initial model, which is evolved with the well–balanced version of the CWENO5 scheme. The logical computational domain is given by the box $\Omega_{\ell}=[-20;20]^{3}$, covered by a uniform grid formed by $80\times 80\times 80$ points. This is also the first test where we have used the re-mapping algorithm for the numerical grid described in Sect. III.2, producing a physical computational domain that extends up to $\sim 300M$ in each direction. In particular, the parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.28$, $b=-8.4$, $c=85.0$, $d=-280.0$, $\xi_{c}=10$.
We also stress that in this test we take full advantage of the new algorithm proposed by [27], which allows to convert from the conserved to the primitive variables even in a $\rho=0$ atmosphere in the exterior of the star.
Fig. 8 reports the results of this set–up by showing the one dimensional profiles of various representative quantities at the final time $t=1000$, compared to the exact solution. The small deviations of the velocity and of the scalar $K$ from the equilibrium values are precisely due to the initial perturbation introduced. Fig. 9, on the other hand, shows the evolution of the Einstein constraints and of the central density.
@@ IV.6 The equilibrium TOV star
Refer to caption Refer to caption
Figure 8: Solution for the TOV star obtained with the Well-Balanced CWENO5 scheme.
Refer to caption Refer to caption
Figure 9: Time evolution of the Einstein constraints (left panel) and of the central density (right panel) for the 3D TOV star.
IV.7 Single puncture black hole
As next test case we consider a single puncture black hole, as originally proposed in [105] (the so–called puncture (or trumpet) solution). The initial condition for the conformal metric tensor is simply the identity matrix, i.e. $\tilde{\gamma}_{ij}=\delta_{ij}=\mathbf{I}$, while the conformal factor at the initial time is set to $\Psi=1+M/(2r)$ with $r=\left\|\mathbf{x}-\mathbf{x}_{c}\right\|$. The black hole is centered in the origin $\mathbf{x}_{c}=(0,0,0)$, has unit mass $M=1$ and zero spin. The initial extrinsic curvature and the initial shift are set to zero, i.e. $K_{ij}=0$, $\beta^{i}=0$, while the lapse is initialized with $\alpha=\psi^{-2}$. The (logical) three dimensional computational domain is given by $\Omega_{\ell}=[-13,13]^{3}$ covered by $132^{3}$ gridpoints with grid-stretching activated. The parameters of the polynomial stretching in the grid function (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-176,176]^{3}$. To avoid the singularity in $r=0$, the computational mesh is built in such a way that no grid point coincides with the origin $\mathbf{x}_{c}$, i.e. the grid points are located in the centers of the equidistant control volumes of the logical grid described above, see also Section III.1 for details. The gamma–driver is turned on, i.e. $s=1$, and the damping parameter in the gamma driver is set to $\eta=2$. We furthermore set $\mu=0$. We have solved this problem with the CWENO7 scheme using exponent $r=4$ in Eq. (55), $\lambda_{0}=10^{10}$, $\lambda_{C}=10^{6}$, and $\lambda_{L}=\lambda_{R}=1$ in Eq. (51) and stopping the simulation at $t=1000$. The initial condition is imposed as Dirichlet boundary condition on the boundary of the computational domain $\partial\Omega$.
Fig. 10 provides a first glance of the results, by reporting the contour plots of the lapse in the $xy$ plane (left panel) as well as the time evolution of the Einstein constraints (right panel). Limited to this test, we have taken the opportunity to perform a close comparison among the results obtained with our new first–order formulation (FO-BSSNOK) and the much more popular second–order formulation (SO-BSSNOK). Hence, Fig. 11 presents such a comparison by reporting the 1D profiles of $\alpha$, $\beta^{1}$, $K$ and $\psi$ in the two cases. Although the numerical schemes adopted are different, namely a seventh order CWENO for FO-BSSNOK and a sixth order linear central finite difference for SO-BSSNOK, the various profiles match perfectly well at the final time $t=1000$. Moreover, also the Einstein constraints, compared in the right panel of Fig. 10, manifest essentially the same trend.
Refer to caption Refer to caption
Figure 10: Solution of the single puncture black hole in a 3D calculation at $t=1000$. Left: contour-plots of the lapse over the $xy$ plane. Right: time evolution of the Einstein constraints.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Comparison among the numerical solutions at time $t=1000$ for the single puncture black hole obtained with (i) our new seventh order CWENO finite difference scheme in the first order FO-BSSNOK system and (ii) a classical sixth order central finite difference scheme applied to the standard second order formulation of BSSNOK (SO-BSSNOK). 1D cuts along the $x_{1}$ axis are shown: lapse $\alpha$ (top left), shift vector component $\beta^{1}$ (top right), trace of the extrinsic curvature $K$ (bottom left) and conformal factor $\psi$ (bottom right).
@@ IV.8 Head on collision of two puncture black holes
Here and in the following Section, we use the procedure initially proposed by [105, 106] to create the initial conditions corresponding to two puncture black holes without excision. The initial metric and the lapse are provided by the TWOPUNCTURES initial data code of [106]. As a first set up, we consider two non–spinning black holes, at rest with respect to each other, and at an initial distance $d=2\,M$. This is the so–called head–on collision. The gamma–driver is necessarily turned on, i.e. $s=1$, again with damping parameter $\eta=2$ and $\mu=0$. We have evolved this configuration with the CWENO7 scheme, over a logical computational domain $\Omega_{\ell}=[-15,15]^{3}$ covered by $132^{3}$ gridpoints, and grid-stretching activated. The parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-335,335]^{3}$. Also in this test case the initial condition is imposed as Dirichlet boundary condition on the boundary of the computational domain $\partial\Omega$.
We have stopped the evolution at $t=1000$, reporting the results in Fig. 12, where we show the contour plots of the scalar $\psi=e^{\phi}$ over the $z=0$ plane. The two black holes merge as expected and at time $t=20$ a single black hole is already formed. The corresponding evolution of the Einstein constraints is instead visible in the left panel of Fig. 14. The constraints slightly increase in time, but only in a very mild form.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12: Head on collision of two black holes obtained with CWENO7.
@@ IV.9 Inspiralling merger of two black holes
Numerical investigations of inspiralling black hole mergers started well before the detection of gravitational waves, and a vast literature exists about their behaviour (see, among the other, [93, 107, 108]). Using again the TWOPUNCTURES library of [106], we place the two black holes at a distance $d=4\,M$ among each other, zero individual spins, opposite linear momenta along the $y$ direction, i.e., $p_{1}=(0,0.45,0)$, $p_{2}=(0,-0.45,0)$, and masses of the two black holes $m_{1}=m_{2}=1$. The gamma–driver is again turned on, i.e. $s=1$, with $\eta=2$ and $\mu=0$. In these conditions we have evolved the system with the CWENO7 scheme over a logical computational domain $\Omega_{\ell}=[-15,15]^{3}$ covered by $132^{3}$ gridpoints, and grid-stretching activated. The parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-335,335]^{3}$. We use the initial condition as Dirichlet boundary condition on the boundary of the computational domain $\partial\Omega$.
We have stopped the evolution at $t=1000$, reporting the results in Fig. 13, where we show the contour plots of the scalar $\psi=e^{\phi}$ over the $z=0$ plane. The two black holes merge after a brief inspiral phase and at time $t=40$ a single black hole is already formed. The corresponding evolution of the Einstein constraints is instead visible in the right panel of Fig. 14.
@@
Figure 14: Time evolution of the Einstein constraints. Left: Head on collision of Fig. 12. Right: Inspiralling merger of Fig. 13
@@ V Conclusions
The second–order BSSNOK formulation of the Einstein equations represents a sort of gold standard in NR, but it also contains an unpleasant feature, in so far it requires quite different numerical schemes for the solution of the Einstein sector and of the matter subsystem. More specifically, writing the Einstein equations as a second–order PDE system prevents from applying to them the best numerical schemes developed over the last decades for hyperbolic systems of conservation laws, which have been developed having first–order systems in mind. In this paper we propose two fundamental advancements in this respect:
@@
Our new implementation can successfully solve all the standard tests of numerical relativity [92], including the long term evolution of a black hole binary system. We believe that the NR community can benefit from this work, both regarding the formulation of the Einstein equations, and the newly proposed numerical scheme, which, being a finite difference method, maintains all the advantages with respect to more expensive DG or finite volume schemes, and yet it can reach any desired high order of accuracy.

### v2

@@ IV.1 Linearized gravitational wave
Figure 1: Linearized gravitational wave test solved with different version of the CWENO scheme. Left panel: $\tilde{A}_{zz}$ component of the extrinsic curvature at the final time, compared to the exact solution. Right panel: time evolution of the normalized Einstein constraints.
@@ IV.2 The robust stability test
Fig. 2 shows the results of our calculations, by reporting the evolution of the normalized Einstein constraints in a battery of tests where the CWENO5 version of the scheme is adopted. These numbers and the corresponding results are certainly satisfactory, even if not of the same quality as those reported in Fig. 4 of [28], where a DG scheme was used in combination with the Z4 formulation of the Einstein equations  [101, 18].
@@ IV.3 The gauge wave
We have first run this test with a small wave amplitude $A=0.01$ over a rectangular domain of size $\Omega=[-0.5,0.5]\times[-0.02,0.02]$ using a CWENO7 scheme. The grid is composed of $256\times 4$ elements, uniformly distributed, and the final time is $t=1000$. As expected, the gauge wave test manifests several pathologies when evolved within the BSSNOK formulation over a long timescale, independently of the numerical scheme adopted. The problem is well known in the literature and has been documented by several authors over the years (see, among others [99, 103, 19]). On the contrary, we recall that this test can be solved successfully within the Z4 or the GH formulations [28, 104, 31]. We have confirmed this behaviour, which is clearly in-printed in the exponential growth of the Einstein constraints, as reported in the left panel of Figure 3. In the right panel we can appreciate the mismatch of the final solution at $t=1000$ with respect to the initial conditions, taking the metric function $\tilde{\gamma}_{xx}$ as a representative quantity.
In spite of these deficiencies, it is still possible to perform a numerical convergence analysis based on this test case, provided the final time is short enough in such a way that the pathologies highlighted above do not have the time to spoil the numerical properties of the scheme. Tab. 2 contains the result of this analysis for a gauge wave with $A=0.1$ at the final time $t=1$ and it essentially confirms the nominal orders of convergence of the scheme.
@@ IV.4 Special relativistic Riemann Problems in the Cowling approximation
All together, these calculation prove the ability of the numerical scheme in treating strong discontinuities in the relativistic regime.
@@ IV.5 Michel accretion
We have solved this test on a computational domain $(r,\theta)\in[0.5;10]\times[0+\epsilon;\pi-\epsilon]$, with $\epsilon=0.05$ and covered by a uniform grid. The relevant parameters are chosen as $r_{c}=8$ and $\rho_{c}=1/16$ like in [110], while an ideal gas with adiabatic index $\gamma=5/3$ is assumed. No well–balancing is used for this test, while analytic boundary conditions are used both at the inner and at the outer radii. Fig. 7 reports the absolute errors (with respect to the exact solution) computed at $t=1000$ while increasing the number $N_{r}$ of radial gridpoints for the CWENO5 scheme, thus showing mesh convergence. In addition, Tab. 4 shows the high order of convergence in a series of companion runs.
Figure 7: Spherical accretion of matter onto a Schwarzschild black hole using Kerr-Schild coordinates. Absolute error at different resolutions for the CWENO5 scheme.
Table 4: Numerical convergence results for the spherical accretion of matter onto a Schwarzschild black hole at a final time $t=10$. The $L^{2}$ errors and the corresponding order of convergence are reported for the variable $\rho$.
$N_{r}$ ${L^{2}}$ error $\rho$ $\mathcal{O}(\rho)$ $N_{r}$ ${L^{2}}$ error $\rho$ $\mathcal{O}(\rho)$
CWENO5 CWENO7
$200$ 4.9639E-05 $200$ 2.7232E-07
$300$ 6.6499E-06 4.96 $300$ 1.9132E-08 6.55
$400$ 1.5811E-06 4.99 $400$ 3.0359E-09 6.40
$500$ 5.1810E-07 5.00 $500$ 7.6260E-10 6.19
IV.6 The equilibrium TOV star
Another fundamental test in numerical general relativity consists of keeping an equilibrium star model stationary over long timescales, providing the first case in our sample of tests where the full Einstein-Euler equations are addressed. Hence, we have first solved the standard Tolman–Oppenheimer–Volkoff (TOV) system [111, 112, 80, 113] for a simple polytropic gas obeying $p=K\,\rho^{\gamma}$.
After choosing the parameters as in [114], namely a central rest mass density $\rho_{c}=1.28\times 10^{-3}$, $K=100$ and $\gamma=2$, we have integrated the ODEs of the TOV system by a Discontinuous Galerkin solver like in [28], obtaining a total mass $M=1.4\,M_{\odot}$ and a radius $R=14.15\,{\textrm{km}}$. A small perturbation in the energy density is then added to the initial model, which is evolved with the CWENO7 version of the scheme. The logical computational domain is given by the box $\Omega_{\ell}=[-15;15]^{3}$, covered by a uniform grid formed by $100\times 100\times 100$ points. This is also the first test where we have used the re-mapping algorithm for the numerical grid described in Sect. III.2, producing a physical computational domain that extends up to $\sim 300M$ in each direction. In particular, the parameters of the polynomial stretching in the grid function of Eq. (72) are $a=2.28$, $b=-68.4$, $c=685$, $d=-2280$, $\xi_{c}=10$.
[…]
@@ IV.6 The equilibrium TOV star
Figure 8: Time evolution of the Einstein constraints (left panel) and of the central density (right panel) for the 3D TOV star solved with the CWENO7 scheme. A comparison among the Well Balanced (WB) scheme and the non Well Balanced (no-WB) is also visible.
IV.7 Single puncture black hole
As next test case we consider a single puncture black hole, as originally proposed in [117] (the so–called puncture (or trumpet) solution). The initial condition for the conformal metric tensor is simply the identity matrix, i.e. $\tilde{\gamma}_{ij}=\delta_{ij}=\mathbf{I}$, while the conformal factor at the initial time is set to $\Psi=1+M/(2r)$ with $r=\left\|\mathbf{x}-\mathbf{x}_{c}\right\|$. The black hole is centered in the origin $\mathbf{x}_{c}=(0,0,0)$, has unit mass $M=1$ and zero spin. The initial extrinsic curvature and the initial shift are set to zero, i.e. $K_{ij}=0$, $\beta^{i}=0$, while the lapse is initialized with $\alpha=\psi^{-2}$. The (logical) three dimensional computational domain is given by $\Omega_{\ell}=[-13,13]^{3}$ covered by $132^{3}$ gridpoints with grid-stretching activated. The parameters of the polynomial stretching in the grid function (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-176,176]^{3}$. To avoid the singularity in $r=0$, the computational mesh is built in such a way that no grid point coincides with the origin $\mathbf{x}_{c}$, i.e. the grid points are located in the centers of the equidistant control volumes of the logical grid described above, see also Section III.1 for details. The gamma–driver is turned on, i.e. $s=1$, and the damping parameter in the gamma driver is set to $\eta=2$. We furthermore set $\mu=0$. We have solved this problem with the CWENO7 scheme using exponent $r=4$ in Eq. (55), $\lambda_{0}=10^{10}$, $\lambda_{C}=10^{6}$, and $\lambda_{L}=\lambda_{R}=1$ in Eq. (51) and stopping the simulation at $t=1000$. In this test we impose non-reflecting Sommerfeld boundary conditions at the outer border of the domain.
The left panel of Fig. 9 provides a first glance of the results, by reporting the time evolution of the Einstein constraints in a series of tests aimed at showing mesh convergence. In addition, and limited to this test, we have taken the opportunity to perform a close comparison among the results obtained with our new first–order formulation (FO-BSSNOK) and the much more popular second–order formulation (SO-BSSNOK). First of all, in the right panel of Fig. 9 we show the Einstein constraints obtained with the two approaches, which manifest essentially the same trend. Furthermore, Fig. 10 enters the details of such a comparison, by showing the errors of the 1D profiles of $\alpha$, $\beta^{1}$, $K$ and $\psi$ using the two schemes, computed at the final time $t=1000$ with respect to the initial condition. In particular, we compare a seventh order CWENO for FO-BSSNOK and a sixth order linear central finite difference scheme for SO-BSSNOK. We note that both methods give pretty much the same error in the inner region close to the puncture.
Concerning a quantitative CPU time comparison, the numerical scheme based on the new FO-BSSNOK is more expensive when compared to the classical SO-BSSNOK formulation, the amount being $87\%$ and $82\%$ at the fifth and seventh orders, respectively. This is essentially due to the larger amount of variables that are evolved in the FO formulation.
Refer to caption Refer to caption
Figure 9: Single puncture black hole: Time evolution of the normalized Einstein constraints as grid resolution is changed (left panel). Comparison of the normalized Einstein constraints among FO-BSSNOK and SO-BSSNOK (right panel).
@@ IV.7 Single puncture black hole
Figure 10: Comparison among the numerical solutions at time $t=1000$ for the single puncture black hole obtained with (i) our new seventh order CWENO finite difference scheme in the first order FO-BSSNOK system and (ii) a classical sixth order central finite difference scheme applied to the standard second order formulation of BSSNOK (SO-BSSNOK). Absolute errors with respect to the initial conditions (computed along 1D cuts) are shown for a few representative quantities: lapse $\alpha$ (top left), shift vector component $\beta^{1}$ (top right), trace of the extrinsic curvature $K$ (bottom left) and conformal factor $\psi$ (bottom right).
@@ IV.8 Head on collision of two puncture black holes
Here and in the following Section, we use the procedure initially proposed by [117, 118] to create the initial conditions corresponding to two puncture black holes without excision.
The initial metric and the lapse are provided by the TWOPUNCTURES initial data code of [118]. As a first set up, we consider two non–spinning black holes, at rest with respect to each other, and at an initial distance $d=2\,M$. This is the so–called head–on collision. The gamma–driver is necessarily turned on, i.e. $s=1$, again with damping parameter $\eta=2$ and $\mu=0$. We have evolved this configuration with the CWENO7 scheme, over a logical computational domain $\Omega_{\ell}=[-15,15]^{3}$ covered by $150^{3}$ gridpoints, and grid-stretching activated. The parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-335,335]^{3}$. Also in this test we impose non-reflecting Sommerfeld boundary conditions at the outer boundaries of the computational domain.
The two black holes merge as expected and at time $t=20$ a single black hole is already formed, that we subsequently evolved until $t=1000$ in order to show long-time stability of our simulation and robustness of both, the underlying mathematical formulation of the equations as well as of the adopted numerical scheme. In Fig. 11 we show the convergence of the normalized Einstein constraints as resolution is increased. The Hamiltonian constraint slightly increases in time, but only in a very mild form.
Figure 11: Time evolution of the normalized Einstein constraints for the Head on collision of Sect. IV.8.
@@ IV.9 Inspiralling merger of two black holes
Numerical investigations of inspiralling black hole mergers started well before the detection of gravitational waves, and a vast literature exists about their behaviour (see, among the other, [100, 119, 120]). Using again the TWOPUNCTURES library of [118], we place the two black holes at a distance $d=4.0\,M$ among each other, zero individual spins, opposite linear momenta along the $y$ direction, i.e., $p_{1}=(0,0.19243,0)$, $p_{2}=(0,-0.19243,0)$, and bare masses of the two black holes $m_{1}=m_{2}=0.46477$. This is one of the models considered by [121] in which the two black holes perform $\sim 2$ orbits before merging. The gamma–driver is again turned on, i.e. $s=1$, with $\eta=2$ and $\mu=0$. In these conditions we have evolved the system with the CWENO7 scheme over a logical computational domain $\Omega_{\ell}=[-6,6]^{3}$ covered by $300^{3}$ gridpoints, and grid-stretching activated. The parameters of the polynomial stretching in the grid function of Eq. (72) are $a=14.59$, $b=-131.33$, $c=395.0$, $d=-394.0$, $\xi_{c}=3$, producing a physical domain of size $\Omega=[-400,400]^{3}$ and an inner maximum grid resolution of $h=0.02$. We impose again Sommerfeld (non-reflecting) boundary conditions.
The merger takes place around $t\sim 70$, after which we let the system evolve to show the stability of the newly formed black hole. The merging phase is shown in Fig. 12, where we report the contour plots of the scalar $\psi=e^{\phi}$ over the $z=0$ plane. The left panel of Fig. 13 shows the trajectories of the two punctures at two different grid resolutions, while in the right panel we report the evolution of the Einstein constraints.
@@
Figure 13: Inspiralling merger of Sect. IV.9 obtained with the CWENO7 scheme. Left: orbits followed by the black holes on the $z=0$ plane. Right: time evolution of the normalized Einstein constraints.
@@ V Conclusions
The second–order BSSNOK formulation of the Einstein equations is one of the most commonly used formulation in numerical general relativity, but it also contains an unpleasant feature, in so far it requires quite different numerical schemes for the solution of the Einstein sector and of the matter subsystem. More specifically, writing the Einstein equations as a second–order PDE system prevents from applying to them the whole class of modern numerical schemes developed over the last decades developed having first–order systems in mind. In this paper we thus propose two fundamental advancements in this respect:
@@
Our new implementation can successfully solve all the standard tests of numerical relativity [99], including the long term evolution of a black hole binary system.
We believe that the NR community could potentially benefit from this work, both regarding the new first order BSSNOK formulation of the Einstein-Euler equations, as well as the proposed CWENO finite difference scheme for its monolithic numerical discretization.